# Holographic analysis of near-conformal dynamics and light dilaton

### v1

@@ Abstract
We carry out a detailed analysis of the region slightly outside the conformal window of a non-trivial infrared fixed point in a generic bottom-up holographic setup. We focus on models, which study the dynamics of a scalar field, dual to quark degrees of freedom, in a (nearly) AdS geometry. Such models realize the picture expected for vector-like near-conformal theories from Dyson-Schwinger analysis. The analysis covers a toy model, which allows for analytic solutions, and a more general setup as well, which encompass a complete model for the ultraviolet physics. We analyze the conditions for the appearance of a parametrically light scalar state in the spectrum, which can act as a candidate for the Pseudo-Nambu-Goldstone boson arising from breaking of the approximate conformal symmetry. We also present detailed results for the vacuum structure, correlators, and Ward identities in the near-conformal regime.
@@ 1 Introduction
Since the discovery of a light Higgs boson at the LHC in 2012 ATLAS:2012yve ; CMS:2012qbp , there have been tremendous efforts to find a hint of new physics beyond the standard model (BSM). One of the expected features for new physics that are consistent with the current status of the experimental endeavor, indicating large separation of scales in BSM, is near conformality, characterized by a parametrically light bound state, corresponding to a pseudo-Nambu-Goldstone boson (PNGB) of the scale symmetry, spontaneously broken by the hyper-quark condensate.
The approach of explaining the extremely fine-tuned Higgs mass from conformal edge, at which the BSM physics enjoys scale invariance at high energies, turns out to be quite attractive. One reason for this is because it nicely explains why there have not been any hint of new particles around the weak scale, $\Lambda_{ew}\sim 1$ TeV, which has been vigorously explored at current energy frontiers. Another notable characteristic is that if an infrared scale is dynamically generated from the quantum criticality, the modulus of the scale-invariant theory, the so-called dilaton, naturally develops a mass as well. If the dilaton is sufficiently light and weakly interacting, it can be a good candidate for dark matter Choi:2012kx .
An appealing approach for our study is to analyze the near-conformal physics from a strongly coupled theory that can be modeled by using the gauge-gravity duality. There have been several studies of holographic models on gauge theories that exhibit a conformal window, with some of them reporting a parametrically light mode, whereas other examples mention the lack of this. We focus here on models that study the dynamics of a scalar field, dual to quark degrees of freedom, in a (nearly) AdS geometry (see, e.g. Matsuzaki:2012xx ; Arean:2012mq ; Evans:2013vca ; Pomarol:2019aae ; CruzRojas:2023jhw ). Such models realize the picture expected for QCD-like theories from the Dyson-Schwinger analysis Kaplan:2009kr . For other holographic approaches discussing near-conformality and light modes, see, for example, Nunez:2008wi ; Elander:2010wd ; Kutasov:2011fr ; Anguelova:2012mu ; Megias:2014iwa ; Cox:2014zea .
@@
There have also been other approaches like in Evans:2013vca where the authors present a simplified holographic model based on the D3-probe-D7 set-up for chiral symmetry breaking in gauge theory. Here, it is shown in an explicit example that if the gradient of the running of the anomalous dimension falls to zero at the scale where the Breitenlohner-Freedman (BF) bound violation occurs, the theory possesses a techni-dilaton state that is parametrically lighter than the dynamically generated quark mass. In Pomarol:2019aae ; Pomarol:2023xcc the authors provide another holographic model dual to a strongly-coupled theory close to the conformal to non-conformal transition, and find that the dilaton mode is always the lightest resonance, although not parametrically lighter than the others.
@@
More recently in a previous work CruzRojas:2023jhw we have also studied a simple holographic model for gauge theories near the conformal edge. By explicitly analyzing the spectra of the composite states, choosing the parameters of the model such that the back reaction to the bulk geometry is negligible, we found that there is a unique scalar state whose mass is parametrically lighter than all other states. The lightest scalar is identified as the Nambu-Goldstone boson associated with the scale symmetry, by showing that it saturates at low energies the anomalous Ward identity of dilatation currents. There are also examples of holographic top-down models that contain light scalar modes, albeit using rather complex Klebanov–Strassler backgrounds Elander:2017cle ; Elander:2017hyr . Simpler models also show light states, but in phases that are unstable (see, e.g., Elander:2021wkc ). Recently, a top-down model with a light dilaton state at the critical end-point of a line of first order phase transition has also been constructed Elander:2025fpk .
@@
In this article, we focus on studying gauge theories that exhibit dynamics slightly outside the conformal windows of non-trivial infrared (IR) conformal fixed points. That is, if the conformal window is bounded below by a critical value $N_{fc}$, in some range of hyper-quark flavors $N_{f}$, we consider values of $N_{f}$ slightly below $N_{fc}$ to be outside the conformal window. We carry out a detailed analysis of this region in a generic bottom-up holographic setup, which includes both the dynamics of chiral symmetry breaking and the renormalization group (RG) flow. To be specific, the RG flow can be modeled in holography by considering Einstein gravity coupled to a flowing dilaton field $\phi$. Then, as pointed out above, the quark degrees of freedom and the description of chiral symmetry breaking are added in terms of another scalar field $X$, dual to the $\bar{q}q$ operator. The interplay of the fields then gives rise to the picture expected from Dyson-Schwinger analysis, where the dimension of the operator $\bar{q}q$ flows from the free field value of three at the high-energy limit to two near the IR fixed point.
@@
Therefore, our setup is significantly more general than earlier studies in this class of models, which typically neglect the UV RG flow or chiral symmetry breaking or both, and restrict to certain models of the IR dynamics. This allows us to clarify which features of the setup are model independent. Note also that having a complete model of the UV flow enables us to analyze one and two-point functions consistently starting from the holographic dictionary.
@@
  • This approximations mean that we explicitly consider a reasonable generic scenario for the UV region (but perhaps not some tuned exotic cases), i.e., basically everything that can happen in this region in a model independent manner. For the IR, which is clearly model dependent, we impose a simple model which should essentially cover all possibilities.
@@
The rest of the article is organized as follows. In Section 2 we introduce the general setup for the holographic model as well as a brief description of the interplay of the holographic RG flow and the chiral symmetry breaking. We also compute the PNGB dilaton mass and give a necessary condition for a light mode to be present. We also show that the masses of the pions, identified as the (flavored) phases of the field which is dual to the quark bilinear operator from our holographic model satisfy the Gell-Mann-Oakes-Renner relation. In Section 3 we consider a simple toy model, which realizes in a concrete way most of the features of a more generic model and allows for analytic computations to be carried out explicitly. In this toy model we consider the flavor field $X$ in an AdS5 background with constant dilaton potential, we also consider a varying mass of the flavor field to account for the scaling dimension of the bilinear operator, and varying boundary conditions as well. We treat the field $X$ in the probe limit and for the UV completion, we directly glue together solutions of $X$ at the IR and the UV. We analyze the possible outcomes for scale separation to see if it exhibits Miransky scaling, and the possibility of having a parametrically light scalar mode in the spectrum. In this section we also consider the effect of the hyper-quark mass on the scale hierarchy to find an Efimov spiral in the plane of the coefficients for the solution of the flavor field in the UV. In Section 4 we consider a more general case where we replace the cosmological constant in the AdS background of the previous toy model by a non-trivial dilaton potential to accommodate the holographic RG flow of the dilaton that corresponds to the running of the coupling of the field theory. Thus we consider explicitly the effect of proper RG flow toward the conformal IR fixed point from a UV free theory (i.e. “UV completion”) until the dynamical generation of scale. We argue that the general model shows the same features as the simple analytic model of the preceding section. We also study the flavor contributions to the one-point and two-point functions of the Yang-Mills sector.
@@ 2 General setup
Before defining specific models, let us discuss the general features of such bottom-up holographic frameworks for QCD-like theories we consider in this article. In general, we expect that the bulk five dimensional action consists of gravity and matter sectors (closed and open string sectors)
@@
Here the gravity action is interpreted to be dual to the Yang-Mills theory, and the matter action is dual to the quark or flavor sector of the theory  [^1] Here and on, we use a term quark for a hyperquark in strongly coupled gauge theories of near conformal dynamics. . In particular, the field $X^{ij}$ is dual to the quark bilinear operator $\bar{q}^{j}q^{i}$ where $i,j$ are the flavor indices. The matter action can include other fields such as the gauge fields but these will not be important to describe the background or fluctuations in the scalar sector, where light modes are potentially expected.
@@ 2.1 Holographic description of the RG flow
We first remind (without complete derivations) what the standard picture of the RG flow for chiral symmetry breaking is and also what Miransky scaling near the conformal transition is. Here we consider the case of zero quark mass. (In the rest of the article, we will also discuss flows with finite quark mass.) As a function of the number of flavors (which we treat as a continuous parameter), the expected phases are the following:
@@
    $N_{f}<N_{fc}$ and $N_{fc}-N_{f}\ll 1$ (walking regime with Miransky scaling). This is the regime that we are mostly interested in as the theory is very close to the conformal window, having a very small $\beta$ function in the IR. There will be two scales, which are widely separated, and we denote the corresponding coordinate values by $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}(\gg r_{\mathrm{UV}})$, and they will be related through the Miransky scaling law. There are three regimes for the solution
@@
    • The IR regime, $r\gtrsim r_{\mathrm{IR}}$. This is the model dependent region where the dilaton and $X$ are nontrivially coupled. Chiral symmetry breaking causes the flow of the dilaton to deviate from the fixed point value. The solutions depend on the chosen IR dynamics, such as the dilaton potential, and/or IR boundary conditions.
@@
where $X_{0}$ and $\alpha$ are constants. One can argue that smoothness of the complete solution at $r\approx r_{\mathrm{UV}}$ and $r\approx r_{\mathrm{IR}}$ requires that these points are close to the nodes of the tachyon. The stability of the solution requires that there are no additional nodes between $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}$; this is the case because in this region of small $X$ field its fluctuations would have essentially the same solution, i.e., a node in the wave function at zero mass, which indicates an instability. The scaling solution is therefore found if $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}$ are set by consecutive nodes. This gives
@@ 2.2 Holographic computation of the dilaton mass
One of our primary concerns is the holographic computation of the mass of the PNGB (Pseudo-Nambu-Goldstone Boson) dilaton arising from the approximate conformal symmetry in the walking regime. In our previous work CruzRojas:2023jhw , it has been noticed that a parametrically light PNGB dilaton, compared to the IR conformal symmetry breaking scale, can be analytically obtained assuming small backreaction from the flavor field $X$. Here we summarize the argument for this and improve the previous discussion. Note that the potential light mode is found in the flavor sector, i.e., it is a meson mode rather than a glueball. In the holographic context this is natural as it is the flavor field $X$ that controls the scale separation and walking, as we discussed above. There is no analogous mechanism in the gluon sector. This is in agreement with several earlier holographic studies, where mesons are typically the lightest states even if they were not parametrically light (see, e.g., Kutasov:2011fr ; Arean:2013tja ; CruzRojas:2023jhw ). Also lattice analyses of near-conformal theories have found signs of such light states Brower:2015owo ; LatKMI:2016xxi ; LatticeStrongDynamics:2018hun .
@@ 2.3 Pions and the Gell-Mann-Oakes-Renner relation
Let us then discuss the masses of the pions and in particular show that they satisfy the Gell-Mann-Oakes-Renner relation at small quark mass. It is well known that holographic models with spontaneously broken chiral symmetry produce this relation with weak assumptions Sakai:2004cn ; Bergman:2007pm . Here we sketch how the analysis works in our case.
@@
and $F^{(L/R)}$ are the field strength tensors of the gauge fields $A_{L/R}$. Note that this is the unique gauge-invariant choice of action up to second order in the fields.
The dependence on the holographic coordinate in the functions $m_{X}$ and $w$ in (2.3) is expected to arise through additional scalar fields. However, it turns out that for the analysis of the pion modes it is enough to consider the fluctuations near the boundary, $r\to 0$. In this region as above we expect that $m_{X}^{2}\approx 3/\ell^{2}$, fixed by the dimension of the dual operator $\bar{\psi}\psi$, while $w$ tends to a constant which value is not relevant for the pion mode. Moreover, the geometry takes the AdS5 form.
@@
where $\hat{\pi}^{a}(r,x)=\pi^{a}(r)e^{-ix_{\nu}q^{\nu}}$ and $X$ is proportional to unit matrix in flavor space. Moreover, we write
@@
Pion fields couple to the longitudinal terms in the axial currents $J_{\mu 5}^{a}=\bar{\psi}\gamma_{\mu}\gamma_{5}t^{a}\psi$. By its definition, the pion decay constant is found as the residue of the longitudinal correlator:
@@
$m_{\pi}^{2}f_{\pi}^{2}=\frac{1}{\kappa^{2}N_{c}}\lim_{r\to 0}\left(\frac{w\ell a_{0}^{\prime}(r)}{r}\right)^{2}\ .$ (28)
@@
$\frac{w\ell a_{0}^{\prime}(r)}{r}=a_{c}+{\cal O}\left(r^{2}\right)\ .$ (30)
@@
Inserting the obtained value of $a_{c}$ in (28) gives the Gell-Mann-Oakes-Renner relation
$m_{\pi}^{2}f_{\pi}^{2}\approx\frac{2}{\kappa^{2}N_{c}}\ell^{3}m_{q}\sigma\ .$ (32)

### v2

@@ Abstract
We carry out a detailed analysis of the region slightly outside the conformal window of a non-trivial infrared fixed point in a generic bottom-up holographic setup. We focus on models, which study the dynamics of a scalar field, dual to quark degrees of freedom, in a (nearly) AdS geometry. Such models realize the picture expected for vector-like near-conformal theories from Dyson-Schwinger analysis. The analysis covers a toy model, which allows for analytic solutions, and a more general setup as well, which encompasses a complete model for the ultraviolet physics. We analyze the conditions for the appearance of a parametrically light scalar state in the spectrum, which can act as a candidate for the Pseudo-Nambu-Goldstone boson arising from the breaking of the approximate conformal symmetry. We also present detailed results for the vacuum structure, correlators, and Ward identities in the near-conformal regime.
@@ 1 Introduction
Since the discovery of a light Higgs boson at the LHC in 2012 ATLAS:2012yve ; CMS:2012qbp , there have been tremendous efforts to find a hint of new physics beyond the standard model (BSM). One of the expected features for new physics that are consistent with the current status of the experimental endeavors, indicating large separation of scales in BSM, is near conformality, characterized by a parametrically light bound state, corresponding to a pseudo-Nambu-Goldstone boson (PNGB) of the scale symmetry, spontaneously broken by the hyper-quark condensate.
The approach of explaining the extremely fine-tuned Higgs mass from the conformal edge, at which the BSM physics enjoys scale invariance at high energies, turns out to be quite attractive. One reason for this is because it nicely explains why there have not been any hint of new particles around the weak scale, $\Lambda_{ew}\sim 1$ TeV, which has been vigorously explored at current energy frontiers. Another notable characteristic is that if an infrared scale is dynamically generated from the quantum criticality, the modulus of the scale-invariant theory, the so-called dilaton, naturally develops a mass as well. If the dilaton is sufficiently light and weakly interacting, it can be a good candidate for dark matter Choi:2012kx ; Hong:2017smd .
An appealing approach for our study is to analyze the near-conformal physics from a strongly coupled theory that can be modeled by using the gauge/gravity duality. There have been several studies of holographic models on gauge theories that exhibit a conformal window, with some of them reporting a parametrically light mode, whereas other examples mention the lack of this. We focus here on models that study the dynamics of a scalar field, dual to quark degrees of freedom, in a (nearly) AdS geometry (see, e.g. Matsuzaki:2012xx ; Arean:2012mq ; Evans:2013vca ; Erdmenger:2020lvq ; Erdmenger:2020flu ; Pomarol:2019aae ; CruzRojas:2023jhw ). Such models realize the picture expected for QCD-like theories from the Dyson-Schwinger analysis Kaplan:2009kr . For other holographic approaches discussing near-conformality and light modes, see, for example, Nunez:2008wi ; Elander:2010wd ; Kutasov:2011fr ; Anguelova:2012mu ; Megias:2014iwa ; Cox:2014zea .
@@
There have also been other approaches like in Evans:2013vca ; Erdmenger:2020lvq ; Erdmenger:2020flu where the authors present a simplified holographic model based on the D3-probe-D7 set-up for chiral symmetry breaking in gauge theory. Here, it is shown in an explicit example that if the gradient of the running of the anomalous dimension falls to zero at the scale where the Breitenlohner-Freedman (BF) bound violation occurs, the theory possesses a techni-dilaton state that is parametrically lighter than the dynamically generated quark mass. In Pomarol:2019aae ; Pomarol:2023xcc the authors provide another holographic model dual to a strongly-coupled theory close to the conformal to non-conformal transition, and find that the dilaton mode is always the lightest resonance, although not parametrically lighter than the others.
@@
More recently in a previous work CruzRojas:2023jhw we have also studied a simple holographic model for gauge theories near the conformal edge. By explicitly analyzing the spectra of the composite states, choosing the parameters of the model such that the back reaction to the bulk geometry is negligible, we found that there is a unique scalar state whose mass is parametrically lighter than all other states. The lightest scalar is identified as the Nambu-Goldstone boson associated with the scale symmetry, by showing that it saturates at low energies the anomalous Ward identity of dilatation currents. There are also examples of holographic top-down models that contain light scalar modes, albeit using rather complex Klebanov–Strassler backgrounds Elander:2017cle ; Elander:2017hyr . Simpler models also show light states, but in phases that are unstable (see, e.g., Elander:2021wkc ). Recently, a top-down model with a light dilaton state at the critical end-point of a line of first-order phase transition has also been constructed Elander:2025fpk .
@@
In this article, we focus on studying gauge theories that exhibit dynamics slightly outside the conformal windows of non-trivial infrared (IR) conformal fixed points. That is, if the conformal window is bounded below by a critical value $N_{fc}$, in some range of hyper-quark flavors $N_{f}$, we consider values of $N_{f}$ slightly below $N_{fc}$ to be outside the conformal window. We carry out a detailed analysis of this region in a generic bottom-up holographic setup, which includes both the dynamics of chiral symmetry breaking and the renormalization group (RG) flow. To be specific, the RG flow can be modeled in holography by considering Einstein gravity coupled to a flowing dilaton field $\phi$. Then, as pointed out above, the quark degrees of freedom and the description of chiral symmetry breaking are added in terms of another scalar field $X$, dual to the $\bar{q}q$ operator. The interplay of the fields then gives rise to the picture expected from Dyson-Schwinger analysis, where the dimension of the operator $\bar{q}q$ flows from the free-field value, $\Delta_{\mathrm{UV}}=3$ at the high-energy limit to $\Delta_{\mathrm{IR}}=2$ near the IR fixed point.
@@
Therefore, our setup is significantly more general than earlier studies in this class of models, which typically neglect the UV RG flow or chiral symmetry breaking or both, and restrict to certain models of the IR dynamics. This allows us to clarify which features of the setup are model-independent. Note also that having a complete model of the UV flow enables us to analyze one and two-point functions consistently starting from the holographic dictionary.
@@
  • This approximations mean that we explicitly consider a reasonable generic scenario for the UV region (but perhaps not some tuned exotic cases), i.e., basically everything that can happen in this region in a model-independent manner. For the IR, which is clearly model dependent, we impose a simple model which should essentially cover all possibilities.
@@
The rest of the article is organized as follows. In Section 2 we introduce the general setup for the holographic model as well as a brief description of the interplay of the holographic RG flow and the chiral symmetry breaking. We also compute the PNGB dilaton mass and give general conditions for a light mode to be present. We also show that the masses of the pions, identified as the (flavored) phases of the field which is dual to the quark bilinear operator from our holographic model satisfy the Gell-Mann-Oakes-Renner relation. In Section 3 we consider a simple toy model, which realizes in a concrete way most of the features of a more generic model and allows for analytic computations to be carried out explicitly. In this toy model we consider the flavor field $X$ in an AdS5 background with constant dilaton potential, we also consider a varying mass of the flavor field to account for the scaling dimension of the bilinear operator, and varying boundary conditions as well. We treat the field $X$ in the probe limit and for the UV completion, we directly glue together solutions of $X$ at the IR and the UV. We analyze the possible outcomes for scale separation to see if it exhibits Miransky scaling, and the possibility of having a parametrically light scalar mode in the spectrum. In this section we also consider the effect of the hyper-quark mass on the scale hierarchy to find an Efimov spiral in the plane of the coefficients for the solution of the flavor field in the UV. In Section 4 we consider a more general case where we replace the cosmological constant in the AdS background of the previous toy model by a non-trivial dilaton potential to accommodate the holographic RG flow of the dilaton that corresponds to the running of the coupling of the field theory. Thus we consider explicitly the effect of proper RG flow toward the conformal IR fixed point from a UV free theory (i.e. “UV completion”) until the dynamical generation of scale. We argue that the general model shows the same features as the simple analytic model of the preceding section. We also study the flavor contributions to the one-point and two-point functions of the Yang-Mills sector.
@@ 2 General setup
Before defining specific models, let us discuss the general features of such bottom-up holographic frameworks for QCD-like theories we consider in this article. In general, we expect that the bulk five-dimensional action consists of gravity and matter sectors (closed- and open-string sectors)
@@
Here the gravity action is interpreted to be dual to the Yang-Mills theory, and the matter action is dual to the quark or flavor sector of the theory  [^1] Here and onwards, we use a term quark for a hyperquark in strongly coupled gauge theories of near conformal dynamics. . In particular, the field $X^{ij}$ is dual to the quark bilinear operator $\bar{q}^{j}q^{i}$ where $i,j$ are the flavor indices. The matter action can include other fields such as the gauge fields but these will not be important to describe the background or fluctuations in the scalar sector, where light modes are potentially expected.
@@ 2.1 Holographic description of the RG flow
We first remind (without complete derivations) what the standard picture of the RG flow for chiral symmetry breaking is and also what Miransky scaling near the conformal transition is. Here we consider the case of zero quark mass; in the rest of the article, we will also discuss flows with finite quark mass. As a function of the number of flavors; which we treat as a continuous parameter, the expected phases are the following:
@@
    $N_{f}<N_{fc}$ and $N_{fc}-N_{f}\ll 1$ (walking regime with Miransky scaling). This is the regime that we are mostly interested in as the theory is very close to the conformal window, having a very small $\beta$ function in the IR. The BF bound is violated, indicating an IR instability, which leads to chiral symmetry breaking, implemented in holography by non-zero $X$ field. There will be two scales, which are widely separated, and we denote the corresponding coordinate values by $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}(\gg r_{\mathrm{UV}})$, and they will be related through the Miransky scaling law. There are three regimes for the solution
@@
    • The IR regime, $r\gtrsim r_{\mathrm{IR}}$. This is the model dependent region where the dilaton and $X$ are non-trivially coupled. Chiral symmetry breaking causes the flow of the dilaton to deviate from the fixed point value. The solutions depend on the chosen IR dynamics, such as the dilaton potential, and/or IR boundary conditions.
@@
where $X_{0}$ and $\alpha$ are constants. One can argue that smoothness of the complete solution at $r\approx r_{\mathrm{UV}}$ and $r\approx r_{\mathrm{IR}}$ requires that these points are close to the nodes of the tachyon. The stability of the solution requires that there are no additional nodes between $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}$; this is the case because in this region where the field $X$ is small, its fluctuations would have essentially the same solution, i.e., a node in the wave function at zero mass, which indicates an instability. The scaling solution is therefore found if $r_{\mathrm{UV}}$ and $r_{\mathrm{IR}}$ are set by consecutive nodes. This gives
@@ 2.2 Holographic computation of the dilaton mass
One of our primary concerns is the holographic computation of the mass of the PNGB (Pseudo-Nambu-Goldstone Boson) dilaton arising from the approximate conformal symmetry in the walking regime. In our previous work CruzRojas:2023jhw , it has been noticed that a parametrically light PNGB dilaton, compared to the IR conformal symmetry breaking scale, can be analytically obtained assuming small backreaction from the flavor field $X$. Here we summarize the argument for this and improve the previous discussion. Note that the potential light mode is found in the flavor sector, i.e., it is a meson mode rather than a glueball. In the holographic context this is natural as it is the flavor field $X$ that controls the scale separation and walking, as we discussed above. There is no analogous mechanism in the gluon sector. [^2] In Sec. 4.3, we will show explicitly that the mixing between the meson mode and the glueball mode is suppressed in the walking regime (i.e. $r_{\mathrm{UV}}\ll r\ll r_{\mathrm{IR}}$). In principle, the mixing can be non-negligible in the deep IR regime ($r\gtrsim r_{\mathrm{IR}}$). However, in our analysis, we assume that the IR physics can be parameterized by a boundary condition near the edge of the walking regime, after cutting off the IR geometry ($r\lesssim r_{\mathrm{IR}}$), so that the mixing is still negligible, and the PNGB dilaton is identified as a meson mode. This is in agreement with several earlier holographic studies, where mesons are typically the lightest states even if they were not parametrically light (see, e.g., Kutasov:2011fr ; Arean:2013tja ; CruzRojas:2023jhw ). Also lattice analyses of near-conformal theories have found signs of such light states Brower:2015owo ; LatKMI:2016xxi ; LatticeStrongDynamics:2018hun .
@@ 2.3 Pions and the Gell-Mann-Oakes-Renner relation
Let us then discuss the masses of the pions and in particular show that they satisfy the Gell-Mann-Oakes-Renner relation at small quark mass. It is well known that holographic models with spontaneously broken chiral symmetry produce this relation with weak assumptions Evans:2004ia ; Sakai:2004cn ; Bergman:2007pm . Here we sketch how the analysis works in our case.
@@
and $F^{(L/R)}$ are the field strength tensors of the gauge fields $A_{L/R}$. Note that this is the unique gauge-invariant choice of action up to second-order in the fields.
The dependence on the holographic coordinate in the functions $m_{X}$ and $w$ in (2.3) is expected to arise through additional scalar fields. However, it turns out that for the analysis of the pion modes it is enough to consider the fluctuations near the boundary, $r\to 0$. In this region, as above, we expect that $m_{X}^{2}\approx 3/\ell^{2}$, fixed by the dimension of the dual operator $\bar{q}q$, while $w$ tends to a constant whose value is not relevant for the pion mode. Moreover, the geometry takes the AdS5 form.
@@
where $\hat{\pi}^{a}(r,x)=\pi^{a}(r)e^{-ix_{\nu}q^{\nu}}$ with $t^{a}\,(a=1,\dots,N_{f}^{2}-1)$ being the generators of $SU(N_{f})$ the fundamental representation and $X$ is proportional to unit matrix in flavor space with flavor-independent quark masses for simplicity. We assume the normalization $\mathrm{Tr}(t^{a}t^{b})=\delta^{ab}/2$. Moreover, we write
@@
Pion fields couple to the longitudinal terms in the axial currents $J_{\mu 5}^{a}=\bar{q}\gamma_{\mu}\gamma_{5}t^{a}q$. By its definition, the pion decay constant is found as the residue of the longitudinal correlator:
@@
$m_{\pi}^{2}f_{\pi}^{2}=\frac{2}{\kappa^{2}N_{c}}\lim_{r\to 0}\left(\frac{\sqrt{w}\ell a_{0}^{\prime}(r)}{r}\right)^{2}\ .$ (28)
@@
$\frac{\sqrt{w}\ell a_{0}^{\prime}(r)}{r}=a_{c}+{\cal O}\left(r^{2}\right)\ .$ (30)
@@
Inserting the obtained value of $a_{c}$ in (28) gives the Gell-Mann-Oakes-Renner relation [^4] Here note that we assume the flavor-independent background $X$ with flavor-independent quark mass $m_{q}$.
$m_{\pi}^{2}f_{\pi}^{2}\approx\frac{4}{\kappa^{2}N_{c}}\ell^{3}m_{q}\sigma\ .$ (32)
The relation of $\sigma$ to the chiral condensate may be computed by following the holographic dictionary (see Appendix E):
$-\frac{2N_{f}\ell^{3}}{\kappa^{2}N_{c}}\sigma=\sum_{i=1}^{N_{f}}\langle\bar{q}_{i}q_{i}\rangle=N_{f}\langle\bar{q}_{i}q_{i}\rangle\ ,$ (33)
where in the last expression there is no sum over the flavor indices. Inserting this in (32) gives
$m_{\pi}^{2}f_{\pi}^{2}\approx-2m_{q}\langle\bar{q}_{i}q_{i}\rangle$ (34)
where the quark condensate $\langle\bar{q}_{i}q_{i}\rangle$ is actually flavor-universal.